OpenStudy (anonymous):

The area of a rectangular piece of land is 260 square meters. If the length of the land was 5 meters less and the width was 2 meters more, the shape of the land would be square. Part A: write an equation to find the width (x) of the land would be a square. show the steps of your work. Part B: what is the width of the land in meters? show the steps of your work.

OpenStudy (anonymous):

@jdoe0001

OpenStudy (anonymous):

@iambatman

OpenStudy (anonymous):

@Luigi0210

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@Compassionate

OpenStudy (anonymous):

@robtobey

hero (hero):

The rectangular area: xy = 260 The square area: (x + 2)(y - 5) = 260 Now, the thing about this is, if the area is square, then x = y, so (x + 2)(x - 5) = 260 x^2 - 3x - 10 = 260 x^2 - 3x - 270 = 0 (x - 15)(x + 18) = 0 x - 15 = 0 x + 18 = 0 Remember, only the positive result matters.

OpenStudy (anonymous):

@Hero so the answer for part A would be this equation : (x+2)(y-5)=260 ?

hero (hero):

The useful equation would be (x + 2)(x - 5) = 260, since we can use it to solve for x.

OpenStudy (anonymous):

@hero thanks, but how would I find the width of the land in meters?

hero (hero):

Solve x - 15 = 0

OpenStudy (anonymous):

x=15

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

thanks a lot, i fully understand how it's done now! :)

hero (hero):

@MarthaSpeaks, very well then :)

OpenStudy (lena772):

So width is 15 for b?